OpenStudy (anonymous):

If the equation(a^2+b^2)x^2 - 2b(a+c)x + b^2+c^2 has equal roots then , which of the following is right and how ? a) 2b=a+c b)b^2=ac c)b=2ac/a+c d)b=ac

OpenStudy (kc_kennylau):

a, b, c are constants?

hartnn (hartnn):

must be

OpenStudy (anonymous):

yeah !

hartnn (hartnn):

you know about discriminants ?

OpenStudy (anonymous):

yes . b^2-4ac !

hartnn (hartnn):

\(b^2-4ac\) thingi ?

hartnn (hartnn):

yeah, so just identify a,b,c here...

OpenStudy (kc_kennylau):

In a quadratic equation \(Ax^2+Bx+C=0\), if \(B^2-4AC>0\), it has 2 solutions. if \(B^2-4AC<0\), it has no real solutions. if \(B^2-4AC=0\), it has only 1 solution.

OpenStudy (anonymous):

i have learnt these rules !

OpenStudy (kc_kennylau):

cool :)

OpenStudy (anonymous):

a = (a^2+b^2 b= - 2b(a+c) c= b^2+c^2

OpenStudy (kc_kennylau):

yep :)

hartnn (hartnn):

correction : a quadratic equation CANNOT have only one solution

OpenStudy (kc_kennylau):

but I would use capital letters A, B, C

hartnn (hartnn):

when B^2-4AC = 0 we have 2 EQUAL solutions

OpenStudy (kc_kennylau):

to avoid ambiguity

OpenStudy (anonymous):

ok !

OpenStudy (kc_kennylau):

@saravananr yep @hartnn is right sorry

hartnn (hartnn):

you want to know where those rules came from ?

OpenStudy (anonymous):

that's ok ! it's understandable !

OpenStudy (anonymous):

i need the way to solve it ? i tried by discriminant method ! but couldn't come to the conclusion

hartnn (hartnn):

\(\huge \dfrac{-b \pm \color{blue}{\sqrt{b^2-4ac}}}{2a}\)

OpenStudy (anonymous):

quadratic formula !

OpenStudy (kc_kennylau):

So you need to set up an equation equating the discriminant and 0 :)

OpenStudy (anonymous):

i did so !

hartnn (hartnn):

yes, from the formula think when we will have 2 equal solutions ? when everything after \(\pm\) equals 0

OpenStudy (anonymous):

-b/2a

hartnn (hartnn):

lets pause this problem for some time and try to understand those rules ? you can resume the problem any time u want....

OpenStudy (anonymous):

ok !

hartnn (hartnn):

because those rules are very important

OpenStudy (anonymous):

ok

hartnn (hartnn):

so you got how the rule came for equal solutions ?

OpenStudy (anonymous):

i know ! i have learnt it

hartnn (hartnn):

ohh, all rules ?

OpenStudy (anonymous):

only major ones !

hartnn (hartnn):

i mean, if you have any doubt in those rules, get them clarified first...

OpenStudy (anonymous):

no , i do not have

hartnn (hartnn):

good, then we can resume your problem,

OpenStudy (anonymous):

best

OpenStudy (kc_kennylau):

so now you can set an equation that B^2-4AC=0 :)

OpenStudy (kc_kennylau):

or in other words, B^2=4AC

OpenStudy (anonymous):

i am trying the same

OpenStudy (kc_kennylau):

[-2b(a+c)]^2=4(a^2+b^2)(b^2+c^2)

OpenStudy (anonymous):

yeah

OpenStudy (kc_kennylau):

b^2(a+c)^2=(a^2+b^2)(b^2+c^2)

OpenStudy (anonymous):

if i am not wrong , i have derived 2ab^2c-a^2c^2-b^4

OpenStudy (anonymous):

i meant 2ab^2c-a^2c^2-b^4=0

OpenStudy (kc_kennylau):

\[\Large\begin{array}{rcl} b^2(a+c)^2-(a^2+b^2)(b^2+c^2)&=&0\\ (a^2b^2+2ab^2c+b^2c^2)-(a^2b^2+a^2c^2+b^2c^2+b^4)&=&0\\ 2ab^2c-a^2c^2-b^4&=&0\\ -1\times \color{blue}{b^4}+2ac\times \color{blue}{b^2}-a^2c^2&=&0 \end{array}\]

OpenStudy (kc_kennylau):

Notice that all options have \(b\) as the subject :)

hartnn (hartnn):

ok, what u got was correct

OpenStudy (anonymous):

yeah !

OpenStudy (anonymous):

but what is the next step to get to the right ans ?

hartnn (hartnn):

couldn't you notice that what u got was a perfect square trinomial?

hartnn (hartnn):

like \(A^2+2AB+B^2\)

OpenStudy (kc_kennylau):

@hartnn wow even I didn't notice it until you reminded me lol

hartnn (hartnn):

with -2AB insteead

hartnn (hartnn):

like \(A^2-2AB +B^2 = (A-B)^2\)

OpenStudy (kc_kennylau):

@hartnn I was like wow there comes a new quadratic equation but I didn't know it's perfect :)

OpenStudy (anonymous):

so is it -(ac+b^2)^2 = 0 ?

hartnn (hartnn):

2ab^2c-a^2c^2-b^4=0 (b^4 - 2(b^2)ac +(a^2c^2)) = 0

OpenStudy (anonymous):

oh ! gr8 !

hartnn (hartnn):

no... (b^2 - ac)^2 =0

OpenStudy (kc_kennylau):

\[\Large\begin{array}{rcl} b^2(a+c)^2-(a^2+b^2)(b^2+c^2)&=&0\\ (a^2b^2+2ab^2c+b^2c^2)-(a^2b^2+a^2c^2+b^2c^2+b^4)&=&0\\ 2ab^2c-a^2c^2-b^4&=&0\\ -1\times \color{blue}{b^4}+2ac\times \color{blue}{b^2}-a^2c^2&=&0\\ b^4-2acb^2+a^2c^2&=&0\\ (ac-b^2)^2&=&0\\ ac-b^2&=&0\\ b^2&=&ac\\ b&=&\sqrt{ac} \end{array}\] All credits to @hartnn :D

OpenStudy (anonymous):

got it very well ! Thanks lot to both of u !

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

both of u have helped me immensly . tomorrow is my exam . plz pray for me

hartnn (hartnn):

wish you a very good luck! :)

OpenStudy (anonymous):

thanks

OpenStudy (kc_kennylau):

gd luck :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@kc_kennylau , and @hartnn r u there ?

OpenStudy (anonymous):

i wan to share smth more

hartnn (hartnn):

yes

hartnn (hartnn):

sure, go on

OpenStudy (anonymous):

i am student of class X living in KSA

OpenStudy (anonymous):

i am an indian

OpenStudy (kc_kennylau):

im from hong kong, close to you :D

OpenStudy (anonymous):

I have started an educational community for the students of both class IX and X where a lot many question papers revision notes and lot more are shared by various student from different parts of the world

OpenStudy (anonymous):

meant for CBSE students

OpenStudy (anonymous):

so , plz u can also contribute to the community by posting smth gud and educational

OpenStudy (anonymous):

the link is iisjcommunity.co.nr

OpenStudy (anonymous):

plz join and spread it to someone whom u all know too !

OpenStudy (anonymous):

hope u will do this help ?

OpenStudy (anonymous):

I just joined saravananr !!

OpenStudy (anonymous):

thanks